OpenStudy (anonymous):

Factor the trinomail completely. 5x^2-75x+50

OpenStudy (zehanz):

First, try to see if there are common factors. In this case all three terms can be divided by 5, so 5 is a common factor: \[5x^2-75x+50=5(x^2-15x+10)\]Now use the "sum and product method" (don't know how it's called in English...): Look for two numbers a and b, with a+b = -15 and a*b = 10. If you've found them, you can write the thing as\[5(x+a)(x+b)\]and you're done!